# The Eye as a Window to Systemic Health: A Survey of Retinal Imaging from Classical Techniques to Oculomics

### v1

@@ II-B1 Pioneer Surgical and observational Techniques
The study of the eye dates to ancient civilisations, where early contributions focused on ocular anatomy, diseases, and basic treatments. Cataracts were among the first documented eye diseases, dating back to around 600 BCE. The Sushruta Samhita, an ancient Indian medical text, described a surgical technique called “couching”, where the opaque lens was inserted into the vitreous humour to restore some vision [33]. However, insertion often led to complications, for instance infection and blindness, underscoring the limitations of early surgical techniques. Greek scholars Hippocrates (460–375 BCE) and Galen (129–216 CE) founded anatomical descriptions of the choroid, lens, and optic nerve. Their work provided the foundation for understanding ocular anatomy, but their reliance on animal dissection led to inaccuracies [34]. Islamic scholars like Al-Zahrawi and Ibn al-Haytham improved surgical tools and advanced optical studies, connecting anatomy with light refraction. Ibn al-Haytham’s Book of Optics emphasised the principles of vision and optics, enabling a framework for understanding refractive errors [35].
@@ II-B2 Early Breakthrough in the optical imaging
The development of the ophthalmoscope by Hermann von Helmholtz in 1851 was a transformative milestone. This device enabled clinicians to visualise the retina and optic nerve directly, facilitating diagnosing of conditions such as retinal detachment and optic neuritis [36]. However, the initial iterations of the ophthalmoscope required considerable skill for effective use, limiting its broader adoption in clinical practice. The early twentieth century saw significant progress in tackling prevalent eye diseases such as trachoma and cataracts. Trachoma, a bacterial infection caused by Chlamydia trachomatis, was a leading cause of blindness in low-income regions. Public health campaigns focusing on hygiene, antibiotic use, and surgical interventions significantly reduced its prevalence. However, these efforts were limited to localised outbreaks and lacked global coordination [37].
@@ II-B3 Technological and optical imaging advancement during the Mid-20th Century
The invention of fundus photography in the 1920s enabled static documentation of the retina, providing a foundation for modern retinal disease research. While this innovation allowed for detailed visualisation, early cameras suffered from low resolution and were not widely accessible. Cataract extraction methods improved with the introduction of intracapsular cataract extraction. However, the lack of intraocular lenses meant that visual outcomes were suboptimal, requiring patients to use thick corrective lenses post-surgery.
The mid-20th century saw rapid advancements in imaging and treatment technologies. In the 1960s, fluorescein angiography allowed for dynamic retinal and choroidal vasculature imaging. It became a critical tool for diagnosing retinal vein occlusion and diabetic macular edema (ME). However, the invasive nature of dye injection posed risks to patients with allergies or comorbidities [38]. Laser treatments became the standard for managing proliferative DR (PDR) and retinal detachment. While effective, these treatments were limited to advanced stages of the disease and could not prevent disease onset. Tonometry was standardised as a reliable method for measuring intraocular pressure, enabling earlier detection of glaucoma. However, tonometry alone could not detect optic nerve damage, highlighting the need for complementary diagnostic tools [39, 40].
@@ II-B4 The Emergence of Digital Era and AI in Ophthalmology
The digital era introduced technologies that significantly enhanced the precision and accessibility of ophthalmic diagnostics. Optical coherence tomography (OCT), invented in 1991, provided cross-sectional imaging of retinal layers with micrometre resolution . This technology revolutionized the management of MD , DR , and GL, but early devices were expensive and required significant expertise [41]. Early computational models automated tasks like vessel segmentation and lesion detection. While these methods were groundbreaking, they were limited by small datasets and reliance on handcrafted features. Digital FP enabled large-scale DR screening programs. These initiatives demonstrated the feasibility of integrating technology into public health but faced challenges in reaching rural and underserved populations [32, 11]. The integration of AI into eye disease research began with classical machine learning (ML) algorithms, which laid the foundation for automated analysis of retinal images and clinical data. Early ML models, such as Support Vector Machines (SVM), Random Forests (RF), and k-Nearest Neighbours (k-NN), were used for tasks like lesion detection, disease classification, and retinal vessel segmentation. These models relied on handcrafted features extracted from FP and OCT images, which limited their generalizability and scalability [42, 43, 44, 45] For example, SVMs were employed to classify DR stages based on features, namely microaneurysms (MAs), hemorrhages(HMs), and exudates(EXs). However, these models struggled with the variability in image quality and the complexity of retinal structures. Despite these limitations, classical ML algorithms demonstrated the potential of AI in automating ophthalmic diagnostics and reducing the burden on clinicians.
@@ II-B5 Early Deep learning models in Eyes research
The advent of deep learning, particularly Convolutional Neural Networks (CNNs), marked a paradigm shift in the analysis of ophthalmic imaging. Unlike classical ML, CNNs automatically learn hierarchical features from raw images, eliminating the need for manual feature engineering. This capability made CNNs highly effective for tasks like DR detection, glaucoma diagnosis, and age-related macular degeneration (AMD) classification. One of the earliest breakthroughs in deep learning for ophthalmology was the development of algorithms for the screening of DR using fundus images. Studies demonstrated that CNNs could achieve a good diagnostic accuracy comparable to that of human experts [46, 47]. For example, the InceptionV3 and ResNet architectures were widely adopted for DR classification, achieving high sensitivity and specificity in detecting early-stage disease. Similarly, CNNs were applied to OCT images for the automated detection of retinal pathologies such as ME, choroidal neovascularization, and retinal detachment. These models improved diagnostic accuracy and reduced the time required for image analysis, enabling faster decision-making in clinical settings.
@@ II-B6 The Rise of pre-trained AI Models and Multimodal Data Integration
CNNs revolutionized ophthalmic imaging, but their limitations in capturing long-range dependencies and spatial relationships led to the development of Vision Transformers(ViTs) [48]. Inspired by the success of transformers in natural language processing, ViTs introduced self-attention mechanisms to image analysis, enabling better modelling of global context and spatial relationships within medical and optical imaging. ViTs have been applied to tasks like DR grading, AMD classification, and retinal vessel segmentation, outperforming traditional CNNs in many cases. For example, Swin Transformers and Multiple Instance Learning ViTs (MIL-ViTs) have been used to detect subtle retinal changes associated with systemic diseases like diabetes and hypertension [49].
In addition, integrating multimodal data, which combines imaging, text, and molecular data, or may of the same modalities, has opened new frontiers in oculomics. For instance, AI models can now analyse retinal images alongside EHRS and genomic data to predict disease progression and treatment outcomes. This multimodal approach has been particularly useful in studying complex conditions such as DR, and AMD, where genetic factors play a significant role in disease pathogenesis [50, 51].
The application of AI extends beyond imaging to the realm of genomics and gene editing. Advances in next-generation sequencing and single-cell RNA sequencing have generated vast amounts of genomic and transcriptomic data, which can be analysed using AI to identify disease-associated genes and pathways. For example, AI-driven genomic analysis has identified key genetic variants associated with AMD, DR and glaucoma [52, 53, 54]. These insights have paved the way for gene editing technologies such as CRISPR-Cas9, which hold promise for treating inherited retinal diseases. AI models are also used to optimise gene editing protocols, predict off-target effects, and design personalised therapeutic interventions for individual patients [55].
@@ II-B7 LLMS and Future AI in the OCULOMICS
The emergence of Large Language Models (LLMs) such as GPT-4 has further expanded the scope of AI in healthcare. LLMs can analyse vast amounts of textual data, including scientific literature, clinical notes, and patient records, to extract insights and generate hypotheses [56]. However, in the field of oculomics, LLM is in the early stages and requires improvement; some study results show [57, 58] that knowledge gained by LLM is usually taken by humans. Although LLMS represent a cutting-edge branch of AI with growing potential in oculomics, it is equally important to understand the broader trajectory of AI as it intersects with medical and retinal research. The evolution of AI, from early origins to present-day applications, forms the foundation for modern techniques that now reshape retinal imaging.

### v2

@@ II-B1 Pioneer Surgical and observational Techniques
The study of the eye dates to ancient civilisations, where early contributions focused on ocular anatomy, diseases, and basic treatments. Cataracts were among the first documented eye diseases, dating back to around 600 BCE. The Sushruta Samhita, an ancient Indian medical text, described a surgical technique called “couching”, where the opaque lens was inserted into the vitreous humour to restore some vision Chan (2010). However, insertion often led to complications, for instance infection and blindness, underscoring the limitations of early surgical techniques. Greek scholars Hippocrates (460–375 BCE) and Galen (129–216 CE) founded anatomical descriptions of the choroid, lens, and optic nerve. Their work provided the foundation for understanding ocular anatomy, but their reliance on animal dissection led to inaccuracies Ghadiri (2024). Islamic scholars like Al-Zahrawi and Ibn al-Haytham improved surgical tools and advanced optical studies, connecting anatomy with light refraction. Ibn al-Haytham’s Book of Optics emphasised the principles of vision and optics, enabling a framework for understanding refractive errors Reeves & Taylor (2004).
@@ II-B2 Early Breakthrough in the optical imaging
The development of the ophthalmoscope by Hermann von Helmholtz in 1851 was a transformative milestone. This device enabled clinicians to visualise the retina and optic nerve directly, facilitating diagnosing of conditions such as retinal detachment and optic neuritis McMullen (1917). Although initial iterations of the ophthalmoscope required considerable skill to be effective, this limited its wider adoption in clinical practice. The early twentieth century saw significant progress in tackling prevalent eye diseases such as trachoma and cataracts. Trachoma, a bacterial infection caused by Chlamydia trachomatis, was a leading cause of blindness in low-income regions. Public health campaigns focusing on hygiene, antibiotic use, and surgical interventions significantly reduced its prevalence. Nevertheless, these efforts were limited to localised outbreaks and lacked global coordination Clare et al. (2024).
@@ II-B3 Technological and optical imaging advancement during the Mid-20th Century
The invention of fundus photography in the 1920s enabled static documentation of the retina, providing a foundation for modern research of retinal disease. Although this innovation allowed for detailed visualisation, early cameras suffered from low resolution and were not widely accessible. Cataract extraction methods improved with the introduction of intracapsular cataract extraction. Due to the absence of intraocular lenses, visual outcomes were less than ideal, requiring patients to wear thick corrective lenses after surgery.
The mid-20th century saw rapid advances in imaging and treatment technologies. In the 1960s, fluorescein angiography allowed for dynamic retinal and choroidal vasculature imaging. It became a critical tool for diagnosing retinal vein occlusion and diabetic macular edema. However, the invasive nature of dye injection posed risks to patients with allergies or comorbidities Weiss & Papakostas (2022). Laser treatments became the standard for managing proliferative DR (PDR) and retinal detachment. While effective, these treatments were limited to advanced stages of the disease and could not prevent disease onset. Tonometry was standardised as a reliable method for measuring intraocular pressure, enabling earlier detection of glaucoma. Tonometry alone cannot detect optic nerve damage, highlighting the need for additional diagnostic tools. Moutray et al. (2018); García-Feijoo (n.d.).
@@ II-B4 The Emergence of Digital Era and AI in Ophthalmology
The digital era introduced technologies that significantly enhanced the precision and accessibility of ophthalmic diagnostics. Optical coherence tomography, invented in 1991, provided cross-sectional imaging of retinal layers with micrometre resolution . This technology revolutionized the management of macular degeneration, DR, and glaucoma, but early devices were expensive and required significant expertise Aumann et al. (2019). Early computational models automated tasks like vessel segmentation and lesion detection. While these methods were groundbreaking, they were limited by small datasets and reliance on handcrafted features. Digital fundus photography enabled large-scale diabetic retinopathy screening programs. These initiatives demonstrated the feasibility of integrating technology into public health, but faced challenges in reaching rural and underserved populations Abràmoff et al. (2010); Schmidt-Erfurth et al. (2018). The integration of AI into eye disease research began with classical machine learning algorithms, which laid the foundation for automated analysis of retinal images and clinical data. Early models, such as Support Vector Machines (SVM), Random Forests, and k-Nearest Neighbors, were used for tasks like lesion detection, disease classification, and retinal vessel segmentation. These models relied on hand-crafted features extracted from CFP and OCT images, which limited their generalizability and scalability Remeseiro et al. (2014); Liu et al. (2011); Veiga et al. (2018); GeethaRamani & Balasubramanian (2018) For example, SVMs were employed to classify DR stages based on features, namely microaneurysms (MAs), haemorrhages (HMs), and exudates (EXs). These models faced challenges due to the varying image quality and complex retinal structures. Despite these limitations, classical Machine Learning (ML) algorithms demonstrated the potential of AI in automating ophthalmic diagnostics and reducing the burden on clinicians.
@@ II-B5 Early Deep learning models in Eyes research
The advent of deep learning, particularly Convolutional Neural Networks (CNNs), marked a paradigm shift in the analysis of ophthalmic imaging. Unlike classical ML, CNNs automatically learn hierarchical features from raw images, eliminating the need for manual feature engineering. This capability made CNNs highly effective for tasks like DR detection, glaucoma diagnosis, and age-related macular degeneration classification. One of the earliest breakthroughs in deep learning for ophthalmology was the development of algorithms for the screening of diabetic retinopathy using fundus images. Studies demonstrated that CNNs could achieve a good diagnostic accuracy comparable to that of human experts Gulshan et al. (2016); Ting et al. (2017). For example, InceptionV3 and ResNet architectures were widely adopted for DR classification, achieving high sensitivity and specificity in detecting early-stage disease. Similarly, CNNs were applied to OCT scans for the automated detection of retinal pathologies such as macular edema, choroidal neovascularization, and retinal detachment. These models improved diagnostic accuracy and reduced the time required for image analysis, enabling faster decision-making in clinical settings.
@@ II-B6 The Rise of pre-trained AI Models and Multimodal Data Integration
CNNs revolutionized ophthalmic imaging, but their limitations in capturing long-range dependencies and spatial relationships led to the development of Vision Transformers(ViTs) Dosovitskiy et al. (2021). Inspired by the success of transformers in natural language processing, ViTs introduced self-attention mechanisms to image analysis, enabling better modelling of global context and spatial relationships within medical and optical imaging. ViTs have been applied to tasks like DR grading, age-related macular degeneration classification, and retinal vessel segmentation, outperforming traditional CNNs in many cases. For example, Swin Transformers and Multiple Instance Learning ViTs (MIL-ViTs) have been used to detect subtle retinal changes associated with systemic diseases such as diabetes and hypertension Yu et al. (2021).
In addition, integrating multimodal data, which combines imaging, text, and molecular data, or may of the same modalities, has opened new frontiers in oculomics. For instance, AI models can now analyse retinal images alongside electronic health records and genomic data to predict disease progression and treatment outcomes. This multimodal approach has been particularly useful in studying complex conditions such as diabetic retinopathy and AMD, where genetic factors play an important role in disease pathogenesis Zekavat et al. (2022); Li, Dong, Zhang, Shi, Liu, Sa, Li, Ni & Mei (2024).
The application of AI extends beyond imaging to the realm of genomics and gene editing. Advances in next-generation sequencing and single-cell RNA sequencing have generated vast amounts of genomic and transcriptomic data, which can be analysed using AI to identify disease-associated genes and pathways. For example, AI-driven genomic analysis has identified key genetic variants associated with AMD, DR and glaucoma Fritsche et al. (2016); Gharahkhani et al. (2021); Chen et al. (2023). These insights have paved the way for gene editing technologies such as CRISPR-Cas9, which hold promise for treating inherited retinal diseases. AI models are also used to optimise gene editing protocols, predict off-target effects, and design personalised therapeutic interventions for individual patients Choi et al. (2023).
@@ II-B7 LLMS and Future AI in the OCULOMICS
The emergence of Large Language Models (LLMs) such as GPT-4 has further expanded the scope of AI in healthcare. LLMs can analyse vast amounts of textual data, including scientific literature, clinical notes, and patient records, to extract insights and generate hypotheses He et al. (2025). Although in the field of oculomics, LLM is in the early stages and requires improvement, some results of the study show Antaki et al. (2023); Betzler et al. (2023) that the knowledge gained by LLM is generally taken by humans. Although LLMS represent a cutting-edge branch of AI with growing potential in oculomics, it is equally important to understand the broader trajectory of AI as it intersects with medical and retinal research. The evolution of Artificial Intelligence, from its early origins to present-day applications, forms the foundation for modern techniques that now reshape retinal imaging.